### v3

# Systematic assessment of various universal machine-learning interatomic potentials

Abstract
[Abstract] Machine-learning interatomic potentials have revolutionized materials modeling at the atomic scale. Thanks to these, it is now indeed possible to perform simulations of ab initio quality over very large time and length scales. More recently, various universal machine-learning models have been proposed as an out-of-box approach avoiding the need to train and validate specific potentials for each particular material of interest. In this paper, we review and evaluate four different universal machine-learning interatomic potentials (uMLIPs), all based on graph neural network architectures which have demonstrated transferability from one chemical system to another. The evaluation procedure relies on data both from a recent verification study of density-functional-theory implementations and from the Materials Project. Through this comprehensive evaluation, we aim to provide guidance to materials scientists in selecting suitable models for their specific research problems, offer recommendations for model selection and optimization, and stimulate discussion on potential areas for improvement in current machine-learning methodologies in materials science.
keywords:
universal machine-learning interatomic potentials, verification, machine learning, phonons, formation energy, geometry optimizationkeywords:
universal machine-learning interatomic potentials, verification, machine learning, phonons, formation energy, geometry optimization1 Introduction
Materials simulations at the atomic scale are the backbone of computational materials design and discovery. They rely on the Born Oppenheimer approximation, in which the electrons follow the nuclear motion adiabatically, so that the potential governing the nuclei consists of the electronic energies as a function of the nuclear positions, called "potential energy surface" (PES). Knowledge of the PES allows the identification of stable and metastable atomic configurations from minimum energy search, or the determination of materials properties as thermodynamical averages from molecular dynamics simulations 1. The utility of atom-based materials simulations is thus intimately related to the generation of accurate PESs, which has been possible in the last decades thanks to the advent of density-functional theory (DFT) 2, 3, 4, 5. Nonetheless, this ab initio approach relies on the quantum mechanical solution of the electronic problem whose computational cost scales cubically with system size and can therefore become unaffordable in various significant cases of technological interest such as amorphous solids, interfaces, surfaces, etc. At the other end of the simulation approaches, parametrized approximations of the Born-Oppenheimer PES, known as empirical analytical potentials, or force fields, or "classical” interatomic potentials, have been widely used especially for large-scale materials studies 6. Unfortunately, in particular when complex electron interactions are involved (as in chemical reactions or phase transitions) these approaches cannot usually achieve DFT accuracy, and in addition they have limited applicability and transferability. They cannot therefore be considered as a drop-in replacement for standard ab initio methods. In this context, machine-learning interatomic potentials (MLIPs) have emerged as an in-between solution with computational cost similar to the empirical analytical potentials, but with the promise of achieving an accuracy comparable to DFT hence enabling accurate simulations over very large time and length scales 7. The key difference with respect to the empirical potentials is that the interatomic potentials are now directly obtained via a highly non linear fit of a set of input/target data (in general, at the DFT accuracy), without any a priori assumption on their analytical form 8. In the original formulation, building a MLIP consists of generating a dataset of atomic configurations for the specific material under study, and training (and subsequently validating) the MLIP on these data based on the accurate prediction of some target metrics as, e.g., energies, forces, and stresses 8, 9, 10, 11, 12, 13, 14, 15, 16. This process is highly material-dependent, and usually requires a significant human and computational effort 17. Obtaining an accurate description of the PES without the need for costly DFT computations, but also covering all possible chemical and structural spaces, would be the holy grail of MLIPs. In this framework, graph Neural Networks (GNN) have revealed very effective, in particular in message passing and equivariant 15, 18 NN 19. In general, NN-based methods can be particularly useful for generalizability thanks to their property of “learning locally” that makes the resulting potential less material-dependent 8, 9, 20, 21, 22. Indeed, the first “universal” MLIP (uMLIP), the MEGNet 23 model, exploited a graph network architecture. It was trained on 60000 inorganic crystals in their minimum energy configurations from the Materials Project (MP) 24 database, which covers 89 elements of the periodic table and is based on the Perdew-Burke-Enzerhof (PBE) exchange-correlation functional 25. This model could provide the formation energy as well as a number of other properties. In order to predict forces and stresses, various other models were subsequently developed relying on different datasets. In particular, the M3GNet 26 and CHGNet 27 models relying on equivariant graph neural network architectures were both developed using snapshots from DFT relaxations of the MP structures. The publication introducing CHGNet was also the opportunity for releasing the Materials Project Trajectory (MPtrj) dataset 27, including the DFT calculations for more than 1.5 million atomic configurations of inorganic structures. By including magnetic moments in the training properties, CHGNet aimed at a better description of chemical reactions as charged states influence how atoms connect with others through chemical bonds 27. A re-implementation of M3GNet, MatGL, has been built on the Deep Graph Library and on PyTorch in order to improve usability and scalability 28. Recently, the M3GNet-DIRECT model has been released 29 in the MatGL package 28 with 1.3 million structures and 89 elements, aiming to predict more reliably unseen structures thanks to a new sampling scheme 29. The ALIGNN-FF model 30 was developed to model a diverse set of materials with any combination of 89 elements from the periodic table but relying on a different database of inorganic crystals, JARVIS-DFT 31, which is based on the optB88vdW exchange-correlation functional 32. Recently, the MACE-MP-0 33 relying on the MACE architecture 19 was also proposed. It was trained on the MPtrj dataset and showed outstanding performance on an extraordinary range of examples from quantum-chemistry and materials science 33. Two proprietary models relying on very large databases have also been developed: the GNoME 34 model exploiting the NequIP architecture 15 and the PFP model 35, 36 exploiting the TeaNet architecture 37. On the one hand, GNoME was trained on a database obtained from a complex active learning workflow of the original MP data, resulting in a number of inorganic structures 100 times larger than the MPtrj. On the other hand, the PFP was trained on a large dataset which initially included 107 DFT configurations covering 45 elements 35 and was further extended to cover 72 elements at the moment of the publication 36, with an expected rise to 94 including rare-earth elements and actinides. It is worth mentioning that uMLIPs have also been developed specifically for organic molecules 38, 39, 40, and for metal alloys 41. While this rapid growth of uMLIP models promises an extraordinary impact in the materials science community, nevertheless benchmark efforts are crucially needed to assess and control their performance and effective usefulness 42, 43, 44.
In this paper, we conduct a comprehensive review and evaluation of four different GNN-based uMLIPs: M3GNet-DIRECT 29, CHGNet 27, MACE-MP-0 33, and ALIGNN-FF 30. They have demonstrated the possibility of universal interatomic potentials that may not require retraining for new applications. The evaluation uses three different datasets: for the equation of state, we use the set of theoretical structures employed in Ref. 45, for the phonon calculations, we use the crystalline structures considered in Ref. 46 that have been relaxed with Abinit 47, 48, norm-conserving pseudopotentials 49, 50, and the PBEsol exchange-correlation functional 51 while for all the other tests, we use the VASP-relaxed structures from the Materials Project (MP) 24, 52.
The paper is organized as follows. After a brief overview of the techniques employed in the generation of the different uMLIPs examined in this work (Sec. 2.1), we test and discuss the quality and transferability of the chosen uMLIPs in different types of calculations, and using the three different datasets as explained above. The tests include the calculation of the equation of state (Sec. 2.2), the evaluation of formation energies and optimization of structural parameters (Sec. 2.3), and the calculation of phonon bands (Sec. 2.4). The conclusions are provided in Sec. 3 and the details of the methods employed are given in Sec. 4.
2 Discussion
2.1 Overview of the uMLIP models
In what follows, we give a brief overview of the basic ingredients underlying
the four GNN-based uMLIPs
examined 27, 29, 30, 33.
Depending on the uMLIPs, they might differ for the specific GNN architecture
employed, for the choice of the training set, for the sampling scheme, or for
the properties included in the training. A comprehensive and detailed
explanation is beyond the scopes of the present work, and we refer the
interested reader to the respective original
papers 27, 29, 30, 33 and
related literature 19, 23, 26, 31, 53, 54, 55, 56.
GNN are reviewed, among others, in
Refs. 18, 57, 58, while
Refs. 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 59 address specifically the construction of MLIP
in materials science.
The pre-trained M3GNet-DIRECT 29 model starts from the Materials 3-body Graph Network (M3GNet) architecture 26, which conjugates the flexibility of typical graph techniques with the physically based many-body features (here 3-body) of traditional interatomic potentials like the Tersoff bond-order potential 60.
The many-body interactions (angles) are aggregated to bonds in standard graph convolution steps to update the bond, atom, and global state (properties) information 26.
The main improvement of M3GNet-DIRECT 29 with respect to the original M3GNet 26 is in the sampling of the materials space. This is done starting from the 128-element vector outputs from the M3GNet 26 model trained on the MP formation energies database 24, to which dimensionality reduction, characteristics-sharing-based clustering, and stratification, are subsequently applied, giving the name of DImensionality-Reduced Encoded Clusters with sTratified (DIRECT) sampling to the approach 29. Static DFT calculations on the so-obtained 1.3 million structures produce the final training set 29. The result is an improved performance with respect to the original M3GNet 26, with a better transferability and prediction of configurations with large energies, forces, and stresses 29.
The Crystal Hamiltonian Graph Neural Network (CHGNet 27) is a model pre-trained on the energies, forces, stresses, and magnetic moments from the MPtrj 27 dataset. The GGA/GGA+U mixing compatibility corrections 61, 62, 63 are applied to the energies 27.
In the CHGNet architecture, the angle information is drawn as a pairwise message passing convolution between bonds (bond graph, where bonds are nodes and edges are angles), on top of the bond information which is drawn as a pairwise convolution between atoms (atom graph, where atoms are nodes and edges are bonds) 27.
To constrain the atom features used to predict energy, forces, and stresses by their charge-state
information, the latter is inferred from the magnetic moments and atomic orbital theory before the last convolution layer 27.
In incorporating the site-specific magnetic moments as the input charge states into the GNN, this model aims to capture the chemical interaction variability across different valence states 27. This should be especially important for transition-metal elements, with variable valence states 64.
The pre-trained MACE-MP-0 model 33 relies on MACE 19, an equivariant message passing neural network potential using messages of an order higher than two body. In combining high body order with message passing, MACE achieves superior accuracy as a result of a trade-off between using more layers and increasing the local body order within a single layer 55.
Using 4-body messages 19, the required number of message passing iterations is reduced to two and the steepness of the learning curve is improved 19.
The high-order equivariant message passing neural network model is combined with the atomic cluster expansion (ACE 13, 65), a framework for deriving an efficient body-ordered symmetric polynomial basis
to represent functions of atomic neighborhoods.
ACE naturally extends to
equivariant features and to include variables beyond geometry, such as charges or magnetic moments 66. MACE-MP-0 is pre-trained on the MPtrj 27 dataset, including a variety of magnetic orders 33. The fitting does not explicitly account for the GGA/GGA+U mixing compatibility corrections 61, 62, 63.
The pre-trained ALIGNN-FF model 30 is the extension to derivative prediction and force field (FF) formalism 30 of the Atomistic LIne Graph Neural Network (ALIGNN) models 56, able to capture two and three body
interactions and to output more than 70 materials properties 30. The ALIGNN 56 models
were implemented using the deep graph library (DGL) 67 which allows efficient line graph construction and neural message passing.
As in the case of CHGNet, the message passing occurs both in an atom graph, where atoms are nodes and interatomic bonds are edges, and in a line graph, where bonds are nodes and bond angles are edges 56.
For the atom graph, nine input node features are assigned to each node, based on the atomic species involved: electronegativity, group
number, covalent radius, valence electrons, first ionization
energy, electron affinity, block and atomic volume.
ALIGNN is a part of the Joint Automated Repository for Various Integrated Simulations (JARVIS) infrastructure 31, a set of databases and tools for materials design. The ALIGNN-FF model was trained on the JARVIS-DFT dataset which contains around 75,000 materials and 4 million energy-force entries,
out of which 307,113 were used in the training 30.
For the sake of simplicity, in what follows we use a shorter nomenclature for three out of the four uMLIPs benchmarked: M3GNet-DIRECT 29 will be called M3GNet, MACE-MP-0 33 will be referred to as MACE, and ALIGNN-FF 30 will be termed ALIGNN.
Some key features of the four uMLIPs are listed in Table 1.
| uMLIP label | Model | Version | Model Size | Data Set | Nelements | Data Size |
| ‘CHGNet’ | CHGNet 27 | 0.3.0 68 | 400K 27 | MP-trj 69 | 89 | 1.58M 27 |
| ‘M3GNet’ | M3GNet-DIRECT 29 | 2021.2.8-DIRECT-PES 70 | 1.1M 43 | MPF 24 + DIRECT 29 | 128 | 1.31M 29 |
| ‘MACE’ | MACE-MP-0 33 | 2023-12-03-mace-128-L1_epoch-199 71 72 | 4.7M 44 | MP-trj 69 | 89 | 1.58M 27 |
| ‘ALIGNN’ | ALIGNN-FF 30 | alignnff_wt10 73 74 | 4.0M 75 | JARVIS-DFT 31 | 89 | 75K 30, 31 |
2.2 Equation of state and comparison with all-electron results
As a first test, we use the protocol for the equation of state (EOS) detailed in Ref. 45, where a high-quality reference dataset of EOS for 960 cubic crystal structures is generated by employing two all-electron (AE) codes. This dataset includes all elements from Z = 1 (hydrogen) to Z = 96 (curium). For each element X, four mono-elemental cubic crystals (unaries structures) are considered, in the face-centered cubic (FCC), body-centered cubic (BCC), simple cubic (SC), and diamond crystal structure, respectively. Besides, for each element, six binary cubic oxides (oxides structures) are included, with chemical formula X2O, XO, X2O3, XO2, X2O5, and XO3, respectively. While in Ref. 45 this dataset is used to gauge the precision and transferability of nine pseudopotential-based ab initio codes, we use it here to assess the four tested uMLIPs.
In Fig. 1, the relative errors on the uMLIPs predictions for the equilibrium volume , the bulk modulus , and its derivative with respect to the pressure obtained from a fit 45 of the EOS, are compared with the analogous errors from selected ab initio methods employed in Ref. 45. FLEUR 76 and WIEN2K 77 are AE codes while VASP 52 implements the projector augmented-wave method (PAW) method 78 with a planewave basis set. The errors reported in Fig. 1 are calculated with respect to the average of the AE methods considered here, i.e., FLEUR 76 and WIEN2K 77.
Furthermore, as in Ref. 45, the EOS computed with two different computational approaches and ( and ) are compared through two metrics and . The first is a renormalized dimensionless version of the metrics introduced in Ref. 79:
where the index runs over the explicit calculations of for the different methods and is the average of over the considered volume range. The second metrics, dependent directly on the physically measurable quantities , , and , captures the relative deviation for each of these three parameters between the two computational approaches and :
where , , and are appropriately chosen weights (see Ref. 45). In Eqs. (1)-(2), is a uMLIP and is the average of the AE methods considered here, i.e., FLEUR 76 and WIEN2K 77. Heatmaps of the periodic table with the values of the comparison metrics and obtained with all the different uMLIPs are reported in Supplemental Material Figs. S1-S4.
It should be noted that most structures in the dataset used in this EOS test are not stable in nature, so that they were not likely included in the dataset used to train the uMLIPs. As a consequence, it is not surprising that the uMLIPs are not able to predict the correct energy versus volume curve for a significant fraction of the systems (Figs. S1-S4). Nevertheless, even for those systems for which a physical EOS is obtained, the precision and transferability is still far from the one that can be achieved with state-of-the-art pseudopotential-based ab initio techniques, as can be observed in Fig. 1. This is a very stringent test for uMLIPs, especially given the effort made in Ref. 45 to improve the precision and the transferability of pre-existent pseudopotential tables used for ab initio calculations. Yet these results suggest that uMLIPs predictions should be taken with some caution and, if possible, validated a posteriori via ab initio calculations, especially if the chemical/physical environment under study is not properly included in the training dataset. The precision of uMLIPs can be improved after retraining the model by including additional ab initio data capturing the chemical/physical configurations under investigation 43, 44. However, despite their relevance, these topics are beyond the scope of the present work and are left for future investigation.
2.3 Structural optimization and formation energies
To test the accuracy of the different uMLIPs further, we prepare a dataset with 19998 materials consisting of unary and binary (with 6903 element combinations) phases in MP. It is worthwhile to highlight that the MP database contains structures relaxed with VASP 52 with the PBE functional 25. The distribution of the chosen dataset on the chemical elements is reported in Fig. 2. We use the different uMLIPs to perform energy calculations both with no structural optimization (we call these “one-shot” calculations) and with structural optimization. The latter are of two types: (i) only the atomic positions are relaxed (we call these “ion-relax” calculations), and (ii) both atomic positions and cell parameters (i.e., lattice parameters and angles) are relaxed (we call these “cell-relax” calculations). For the structural optimizations, in some cases the calculation stops due to errors while building the graph representation (e.g., given that isolated atoms are found in the structure). Calculations for which this problem appears are tagged as problematic. In another non-negligible number of cases, that we tag as unconverged, the relaxation algorithm is not able to reach the stopping criterion before 150 steps. Finally, we tag as converged the calculations where the stopping criterion is met in less than 150 steps. Additional details on the relaxation algorithm are given in Sec. 4.
In Fig 3, we report the number of occurrences of the problematic, unconverged, and converged tags from the different uMLIPs for the two types of structural optimization. When only the atomic positions are relaxed (Fig 3(a)), the fraction of unconverged cases is very limited (about 0.30.4% of the total for all the uMLIPs), while only ALIGNN presents some problematic cases worth to be noted, but still very small (0.2%). The fraction of unconverged calculations increases significantly when both atomic positions and cell parameters are optimized (Fig 3(b)), especially for M3GNet (2.5%) and ALIGNN (4.4%). CHGNet and MACE perform more robustly, with lower fractions of unconverged results (both 1.2%). In the following, we discuss only results from the converged geometry optimizations.
To compare the predictive performance of these uMLIPs on the energetics, we use as a benchmark the formation energy, defined as
where is the total energy per atom of the phase of interest with , the constituent elements, and are the fractions of and , respectively, and and are the lowest possible energies per atom of and , respectively. The ability of a uMLIP in predicting formation energies is evaluated as a difference from the MP values, i.e.,
with and being the formation energy (Eq. (3)) from the MP and calculated with the uMLIP model, respectively.
We choose to work with the one-shot energies as this allows to avoid spurious effects due to the different relaxed geometries predicted by the uMLIPs. It is important to note that all our calculations incorporate energy corrections. For CHGNet and M3GNet, these corrections are already encoded in the uMLIPs, as both models were trained using corrected energies 61, 62, 63. Conversely, for MACE, the corrections are applied as a post-processing step (see Sec. 4 and Refs. 81, 82 for details) since this model was trained on uncorrected entries 33. ALIGNN, on the other hand, was trained using the JARVIS-DFT dataset with OptB88vdW functional 32, and does not use any corrected energies, but rather direct DFT outputs. In this case, we compute Eq.(4) with and without energy corrections, and we select the best-agreement set of results.
Figure 4 shows the distribution of the absolute mean values of on the chemical space for the different uMLIPs, i.e., the average of the absolute value of over all the chemical systems containing a given element. CHGNet and MACE have the best performance over all the periodic table. For what concerns the elements in the transition metal series, CHGNet and MACE perform significantly better than M3GNet and ALIGNN, especially for elements like V, Mo, and W. In the case of CHGNet, this might be related to the inclusion of the magnetic moments in the training process. CHGNet and MACE outperform M3GNet and ALIGNN also for what concerns the calchogens (O, S, Se, Te) and some halogens (F, Cl, I). ALIGNN is also less performant for actinide elements Np and Pu compared to other uMLIPs. In Fig. 5, we report the distribution of (Eq.(4)) for the one-shot calculations using the different uMLIPs, while the performance of each model is summarized in Table 2. MACE shows better performance than the other uMLIPs, with the smallest MAE and RMSE and the highest R2 (0.044, 0.101, and 0.989, respectively). In turn, MACE outperforms the other uMLIPs in the one-shot calculations (Fig. 5 and Table 2), while displaying, together with CHGNet, the highest number of successful structural optimizations (or converged results), i.e., 88.8%, while M3GNet and ALIGNN display 87.5% and 85.6%, respectively (Fig. 3).
To examine the ability of the different uMLIPs to predict the cell geometry (i.e., lattice parameters, angles, and volume) in the cell-relax calculations, we compare the uMLIPs results with the MP values, as done for the formation energy (Eq. (4)). E.g., for the volume we consider the relative difference
where and are the cell volume from MP and from the uMLIP model cell-relax calculations, respectively. The distribution of is reported in Fig. 6. CHGNet and M3GNet outperform the other uMLIPs, with a narrower distribution of . ALIGNN performs the worst, while MACE has an intermediate performance. In Table 3, we report the Mean Absolute Relative Error (MARE) on the predicted lattice parameters, angles, and volumes. We observe that MACE shows similar performance with respect to CHGNet and M3GNet for , , and , , , but its performance on the volume is significantly worse than the other two uMLIPs, as also shown in Fig. 6. One possible origin of this behavior might be the compensation of the errors of the cell parameters in the case of M3GNet and CHGNet, which would not occur in MACE. Another reason could be the presence of more outliers (extreme large deviations) in the case of MACE, as documented from the histograms of the relative volume differences and lattice parameters differences provided in the Supplemental Material.
A similar argument can be applied to the slightly higher MARE on the volume for CHGNet as compared to M3GNet (Table 3). ALIGNN shows larger errors compared to the other uMLIPs (Table 3). To further test the ability of uMLIPs to treat multi-element compounds, we perform ‘one-shot’ and ‘cell-relax’ calculations on 100 randomly chosen quinary materials. There are 4 unconverged cases (4%) for CHGNet and M3GNet, while 2 unconverged cases (2%) for MACE and ALIGNN. This seems to indicate that these uMLIPs can still be used for structural relaxations even for quinaries although additional investigations would be needed to assess the transferability in this part of the chemical space. In the Supplemental Material we report the comparison of formation energies and cell geometries of quinaries with the respective MP reference data, as done for the unary and binary compositions (Figs. 5 and 6). This comparison indicates that MACE still gives the most reliable prediction on formation energies, while CHGNet shows the best performance for cell geometry predictions.
| uMLIP | MAE | RMSE | R2 |
| CHGNet | 0.054 | 0.105 | 0.988 |
| M3GNet | 0.172 | 0.316 | 0.896 |
| MACE | 0.044 | 0.101 | 0.989 |
| ALIGNN | 0.137 | 0.2226 | 0.947 |
| uMLIP | volume | a | b | c | |||
| CHGNet | 3.16 | 2.03 | 2.07 | 2.44 | 0.75 | 0.62 | 1.19 |
| M3GNet | 2.97 | 2.04 | 2.09 | 2.46 | 0.89 | 0.73 | 1.24 |
| MACE | 5.22 | 2.01 | 2.11 | 2.58 | 0.73 | 0.59 | 1.13 |
| ALIGNN | 7.85 | 3.42 | 3.42 | 3.61 | 0.94 | 0.86 | 1.32 |
2.4 Vibrational properties
In this section we analyze the capability of uMLIPs to reproduce the vibrational properties of crystalline materials using accurate ab initio results reported in a previous work 46 as a reference. From the structures in Ref. 46 we select those whose energy above hull is zero both in the MP database and from the uMLIP cell-relax calculations, leading to 101 structures. Phonons are computed using the finite displacement method as implemented in the phonopy package 83, 84. Further details on the protocol employed to compute phonons with uMLIPs are provided in Section 4. By exploiting the finite displacement method to compute phonons, this study indirectly probes the quality of the uMLIPs-calculated forces when atoms are slightly displaced away from the equilibrium positions. From a methodological point of view we note that, already when ab initio engines are used, the accuracy of phonon calculations from the finite displacement method is rather sensitive to the quality of the force calculations in the supercell. Since by construction these forces are less accurate when calculated from uMLIP models than from ab initio methods, it is reasonable to expect some non-negligible discrepancy between uMLIPs and ab initio phonon calculations. Also, as discussed in more detail in Sec. 4, the MLIP architectures at the basis of the universal models considered in this study cannot predict the long-range dipolar contributions to the interatomic force constants, needed to obtain a reliable Fourier phonon interpolation in the case of polar materials 85, 86. The uMLIP-calculated forces should therefore be augmented with the ML electronic dielectric tensor and the Born effective charges, which are not available at present, leading us to augment the uMLIPs forces with ab initio 47, 48 values for these quantities (see Sec. 4 for further details).
For the above reasons, when comparing uMLIPs results with ab initio data, we choose a rather generous metrics, i.e., the MAE between the uMLIPs and the ab initio phonon band structures computed along a high-symmetry -path 87:
In Eq. (6), is the phonon energy in meV, is the branch index, is the number of wavevectors used to sample the -path, and , are the phonon energies computed using the uMLIP model and the density-functional perturbation theory (DFPT) with ABINIT 46, respectively. Table 4 reports the minimum, maximum, and average values of the MAE in the phonon band structures obtained from the different uMLIPs (Eq. (6)), while the MAE distribution is given in panel (a) of Fig. 7.
| uMLIP | MIN_MAE | MAX_MAE | MEAN_MAE |
| CHGNet | 0.82 | 37.34 | 8.12 |
| M3GNet | 0.74 | 40.20 | 10.41 |
| MACE | 0.31 | 17.22 | 3.71 |
| ALIGNN | 5.60 | 75.38 | 29.36 |
From panel (a) of Fig. 7 and Table 4, we observe a non-ideal agreement between uMLIPs-calculated and ab initio-calculated vibrational properties. In addition to the fundamental reasons discussed above, in this particular study a source of discrepancy could come from the fact that Ref. 46 exploits the PBEsol exchange-correlation functional while the uMLIPs have been trained using PBE 25 data with the inclusion of U-corrections 61, 62, 63 for certain systems (in the case of CHGNet and M3GNet) or PBE 25 data without the inclusion of U-correction (in the case of MACE) or optB88vdW 32 data (in the case of ALIGNN). In panel (b) of Fig. 7, we report the phonon band structure for the system with the lowest MAE (0.312 meV, obtained from MACE, Table 4), which is Sr4Br8 (mp-567744). Conversely, in panel (c) of Fig. 7, we report the phonon band structure for the system with the highest MAE (75.375 meV, obtained from ALIGNN, Table 4), which is Be2C (mp-1569). For the latter, the uMLIP predicts vibrational instabilities that are not observed in the ab initio results. However, we also note that even for the lowest-MAE material (panel (b) of Fig. 7), the high-energy region of the phonon band structure is still far from being accurate.
To summarize, our results indicate that presently available uMLIPs can predict ab initio vibrational properties with a typical error of 3.71 meV in the best-case scenario. This value should be considered as a lower bound as the discrepancy is expected to increase if the electronic dielectric tensor and the Born effective charges were predicted with ML techniques. In the opinion of the authors, the predictive behavior of uMLIPs for vibrational properties can be improved by training new uMLIPs with larger weights for the forces loss function but it is also clear that for accurate ML-based predictions in polar materials, one needs uMLIPs capable of inferring the long-range part of the dynamical matrix. All this being said, we believe that uMLIPs represent an efficient and promising approach to perform an initial screening for vibrational and thermodynamic properties, especially in a high-throughput context in which high accuracy is not necessarily needed.
3 Conclusions
We present a systematic assessment of various universal machine-learning interatomic potentials (uMLIPs) by investigating their capability to reproduce ab initio results for several important physical properties such as equation of state, relaxed geometries, formation energies, and vibrational properties of an extensive set of crystalline materials. Among the considered uMLIPs, we find that MACE shows superior accuracy in predicting formation energies and vibrational properties, and CHGNet and M3GNet are outstanding for relaxed geometry predictions. MACE and CHGNet show superior performance on the formation energy prediction along the periodic table, especially when considering systems containing transition metal elements, chalcogens, and halogens (with Br being accurately reproduced only by CHGNet). M3GNet exhibits relatively high errors, leading to a lower R2 value compared to the other uMLIPs in predicting formation energies. Despite this, its ability to accurately predict volume and lattice parameters remains high, and it has intermediate performance in predicting vibrational properties. For what concerns ALIGNN, despite trained with different dataset, it still shows relatively good results when predicting formation energies. However, it also reveals to be the most problematic model when performing geometry optimization at variable cell and for predicting vibrational properties. These results underscore the need for further optimization and training of the currently available uMLIPs to fully exploit the capability of ML techniques across a broader range of applications. The choice of a particular uMLIP for specific applications should take into account an appropriate balance between accuracy and computational efficiency. Future work should aim at enhancing the performance of these potentials further, particularly focusing on areas where current uMLIPs exhibit limitations such as more accurate prediction of forces and stresses or the capability of learning Born effective charges and electronic dielectric tensors that are crucial for the vibrational properties of polar materials. Our work will hopefully pave the way towards a more systematic assessment of uMLIPs in different scenarios and the establishment of a standardized benchmark set that can be used to gauge the precision and transferability of uMLIPs.
4 Methods
The calculations are performed with the AbiPy package 47, 48, more specifically the abiml.py script providing a unified interface that allows one to perform different types of calculations such as structural relaxations, molecular dynamics, or NEB using the algorithms implemented in ASE 88 and different uMLIPs as calculators. The following package versions are used to produce the results reported in this work. Python: 3.11, Pymatgen: 2023.7.17, AbiPy: 0.9.6, Abinit: 9.8.4, CHGNet: 0.3.2, ALIGNN-FF: alignn 2023.10.1, MatGL: 1.0.0, MACE-MP0: mace-torch 0.3.4, and ASE: 3.22.1.
Structural relaxations are performed using the ASE optimizer, employing the BFGS algorithm. To ensure good trade-off between accuracy and computational cost, the stopping criterion fmax is set to 0.1 eV/Å. If the stopping criterion is not met at a maximum number of iterations set to 150, the calculation is considered unconverged. This is justified by the fact that all the initial structures in our dataset are taken from the Materials Project 24, where they have been already relaxed with VASP 52 with the PBE functional 25, and the same structures are supposed to be included in the set used to train the uMLIPs (with the exception of ALIGNN). It should be pointed out that for each composition the total energy obtained from the Materials Project 24 is obtained with energy corrections 61, 62, 63. Therefore, we apply the same correction according to the methods in compatibility.py from the Pymatgen package 81, 82 to MACE since it was trained with Materials Project uncorrected energies.
For the phonon calculations we first perform a structural relaxation with the
uMLIPs from the crystalline structures employed in 46.
The atomic positions are relaxed at fixed cell parameters in order to avoid
spurious effects due to the change of the lattice parameters with respect to the
reference DFT results.
The relaxed configuration is then used to compute vibrational properties using
the finite displacement method implemented in
phonopy 83, 84 with a displacement
of 0.01 Å.
In each calculation, the real-space supercell is matched with the -mesh
employed in 46 to compute the dynamical matrix with the DFPT
part of Abinit 47, 48.
It is noteworthy that the uMLIPs employed in this work lack the capability to
predict the electronic dielectric tensor and the Born
effective charge tensor where is the index of the
atom in the unit cell.
As discussed in 85, 86, these quantities are needed to
model the long-range dipolar contribution to the interatomic force constants in
polar materials.
This term is indeed responsible for the LO-TO splitting as well as for the
non-analytical behaviour of the vibrational spectrum near .
Its correct numerical treatment is therefore crucial to obtain an accurate
Fourier interpolation of the dynamical matrix at arbitrary -points.
For this reason, in all phonopy calculations we use the ab initio values of
and obtained with Abinit to model
long-range interactions.
The ab initio phonon dispersions is obtained by using the anaddb post-processing tool using the DDB files retrieved from the MP database to
compare the different phonon dispersions between uMLIPs and ab initio DFPT
results.
Figure 1 and the heatmaps of the periodic table in the
Supplemental Material have been produced using the ab initio results and the
open source python scripts available in
the acwf-verification-scripts
github repository.
*Author contributions
G.-M. R. designed the research topic and coordinated the whole project. All authors provided the ideas underlying this work, contributed to its development, and discussed the findings reported in the paper. All authors took part to the writing and reviewing of the paper, and to the final approval of its completed version. H. Y. and M. G. executed the calculations presented. M. G. developed and maintained the software infrastructure.
*Acknowledgments
The authors acknowledge useful discussions about the correct use of the different uMLIPs with Yuan Chiang, Kamal Choudhary, Bowen Deng, Tsz Wai Ko, and Shyue Ping Ong. Computational resources have been provided by the supercomputing facilities of the Université catholique de Louvain (CISM/UCL), and the Consortium des Equipements de Calcul Intensif en Fédération Wallonie Bruxelles (CECI). This work is supported by the National Key Research and Development Program of China(2022YFE0141100 and 2023YFB3003005).
*Financial disclosure
None reported.
Conflict of interest
The authors declare no potential conflict of interests.